OpenStudy (anonymous):

which of the following are the solutions to the equation below? Check all that apply. 5x^2-2x+16=4x^2+6x a.-3 b.6 c.-4 d.4 e.18 f.-6

OpenStudy (anonymous):

\[5x^2-2x+16=4x^2+6x \]Move everything to the left side.. \[x^2-8 x+16 =0\]

OpenStudy (anonymous):

Got it from here?\[(x-4)^2= 0\]

OpenStudy (anonymous):

wouldnt it be c nd d

OpenStudy (anonymous):

x-4=0 x+4=0 x=-4 x=4 right

OpenStudy (anonymous):

oh snaps ok

OpenStudy (anonymous):

Nope! Close...

OpenStudy (anonymous):

+4 only huh

OpenStudy (anonymous):

\[(x−4)^2=0\]\[x−4=0\]

OpenStudy (anonymous):

yes